### v1

@@ D.4 Additional Results on Model Fit
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 2.441 2.420
(0.409) (0.551)
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$ 2.993 3.270
(0.229) (0.316)
R-squared 0.28 0.38 0.44 0.51
@@
Notes: This table reports the pass-through of tariffs to quantities at the exporter–importer–product level. Panel A presents reduced-form estimates from the data. Panel B shows corresponding pass-through estimates generated by the model. Columns (2)–(4) interact tariff changes with lagged supplier share ($s_{ijht-1}$) and lagged buyer share ($x_{ijht-1}$). Columns (1) and (2) use baseline fixed effects ($FE_{ht}+FE_{cs}$), while Columns (3) and (4) employ a more stringent specification with product–year, country–year, and buyer–year fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$). Standard errors are clustered at the HS8 product and exporter-country level. Observation counts are rounded to four significant digits per U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
@@
$\widehat{\Delta\ln q_{ijht}}$ 0.349 0.525 0.205 0.311 0.185 0.282 0.420 0.627
@@
$\widehat{\Delta\ln(p\cdot q)_{ijht}}$ 0.176 0.426 0.102 0.250 0.092 0.228 0.212 0.509
@@ D.5.1 Pass-Through Across Samples
Panel A and Panel B in Table D.9 evaluate how tariff pass-through estimates vary under different sample restrictions using baseline fixed effects (product–time, country–sector) and more stringent fixed effects (including buyer-time), respectively. For ease of exposition, the first two columns of Table D.9 replicate the numbers of Columns (1) and (3) of Table 5. These two columns show that the baseline sample yields a pass-through estimate of around 78-85%, closely matched by the model-implied prediction with no statistically significant difference. As the sample is progressively broadened–from including single-buyer matches (Column (3)), to adding related-party transactions, energy goods, and outliers (Column (4)), and finally to the most inclusive specification (Column (5))–estimated pass-through increases steadily, reaching up to 93-95%. This pattern underscores the sensitivity of reduced-form estimates to sample composition and the role of relationship filtering in uncovering pricing patterns consistent with bilateral bargaining.
@@
Dependent variable: Model Baseline +Suppliers + Energy/ + Final
$\Delta\ln p_{ijht}$ w/ RPT/ Goods
$<$ 2 Buyers Outliers
@@
R-squared 0.50 0.31 0.18 0.14 0.13
Observations 249,000 249,000 732,000 945,000 1,768,000
Notes: This table reports tariff pass-through estimates to duty-exclusive prices at the exporter–importer–product level. Panel A uses baseline fixed effects: product–time and country–sector ($FE_{ht}+FE_{cs}$). Panel B uses a more stringent specification: product–time, country–time, and buyer–time ($FE_{ht}+FE_{ct}+FE_{jt}$). Column (1) uses the model-predicted price change as the dependent variable. Column (2) uses the observed price change in the baseline sample and is identical to Column (1) of Table 3. Column (3) adds relationships in which the supplier trades with only one U.S. importer. Column (4) further expands the sample to include relationships that are either related parties, involve energy commodities, or exhibit extreme price levels or changes. Column (5) incorporates consumption goods, thus encompassing all consecutive exporter-importer-product combinations. Standard errors are clustered at the HS8 product and country level. Observation counts are rounded per U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
D.5.2 Pass-Through Using Data Aggregated at the Product-Level
To complement our main analysis, we replicate a standard pass-through specification using data aggregated at the product-country-month level. We construct the data directly from the buyer–supplier–product triplets that form the basis of our firm-level regressions. While much of the literature analyzes monthly price and tariff changes at the product–country level, such approaches reflect both intensive and extensive margin adjustments, including changes in trading partners or the entry and exit of relationships. In contrast, our aggregation focuses exclusively on consecutive transactions between the same buyer and supplier for a given product. This setup isolates price responses within ongoing relationships, capturing what is arguably the most direct expression of tariff pass-through at the micro level.
@@ D.5.2 Pass-Through Using Data Aggregated at the Product-Level
We measure price changes at the HS10–country–month level, using the same subset of buyer–supplier–product links as in the baseline analysis. These are links with at least one transaction in both 2017 and 2018. For each product and country, we construct monthly prices by aggregating trade values and quantities. We then relate monthly price changes to changes in statutory tariffs, controlling for product-month, country-month, and country-sector fixed effects. Standard errors are clustered at the HS8–country level.
Models
Table D.10, reports results separately for all products and for the subset that excludes consumption goods. Within each group, columns reflect increasingly selective samples. Columns (1) and (4) includes buyer–supplier–product pairs observed in two consecutive calendar years. The next specification, Columns (2) and (5), excludes consumption goods and applies additional filters: energy-sector goods are dropped; transactions with extreme price levels and price changes are excluded; and related-party trade is removed. The final sample, Columns (3) and (6), is restricted to suppliers trading with at least two U.S. buyers for the same product in consecutive years.
@@ Results
Across specifications, we find consistent evidence of incomplete tariff pass-through to U.S. import prices. As in our firm-level regressions, the degree of pass-through incompleteness increases as we move to more selective samples, particularly those that condition on firms with multiple trading partners over consecutive years. The estimated effects are generally larger in magnitude than at the match level, suggesting that relationship-level frictions may be amplified when observed in aggregated trade flows. While this exercise remains suggestive, it helps connect the mechanisms explored in the main analysis to pricing patterns in product-level data.
@@
Table D.10: Tariff Pass-Through Using Data Aggregated at the Product-Level
Dependent Variable: $\Delta\ln p_{cht}$
All Products Excl. Consumption Goods
(1) (2) (3) (4) (5) (6)
$\Delta\ln(1+\tau_{cht})$0.115 -0.136 -0.321 -0.072 -0.119 -0.700
(0.078) (0.058) (0.186) (0.136) (0.101) (0.164)
Consecutive obs only Yes Yes Yes Yes Yes Yes
(-) Energy/RPT/Outliers No Yes Yes No Yes Yes
(-) Suppliers with $<$2 buyers No No Yes No No Yes
Observations 800,000 540,000 180,000 510,000 320,000 100,000
R-squared 0.13 0.14 0.22 0.13 0.14 0.22
@@
Notes: This table reports regressions of month-over-month changes in log unit values (FOB, excluding charges) on corresponding changes in statutory tariffs ($\Delta\tau_{hct}$), measured at the HS10–country–month level. The data are aggregated from the firm-level sample used in our main analysis, retaining only buyer–supplier–product triplets with consecutive transactions. Columns (1)-(3) refer to all products; Columns (4)-(6) exclude consumption goods, based on the Broad Economic Categories (BEC) classification. Each column reflects a progressively more restricted sample: Columns (1) and (4) include all consecutive transactions; Columns (2) and (5) drop energy goods, extreme price levels and changes, and related-party trade; and Columns (3) and (6) restrict to suppliers with multiple buyers in consecutive years. All regressions include product–month, country–month, and country–sector fixed effects. Standard errors are clustered at the HS8–country level and reported in brackets. The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
D.5.3 Pass-Through Results Using LFTTD RPT Indicators
As a robustness check, we re-estimate the baseline specifications using the related-party transaction (RPT) indicator provided in the LFTTD. Table D.11 shows that the pass-through estimates remain stable, and the interaction effects with supplier and buyer shares are qualitatively similar to those reported in Table 3.
Table D.11: Pass-Through and Relationship Heterogeneity–Alternative Related Party Trade
Dependent variable: $\Delta\ln p_{ijht}$
(1) (2) (3) (4)
$\Delta\ln(1+\tau_{cht})$0.163 -0.087 -0.304 -0.253
(0.096) (0.100) (0.111) (0.114)
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.043 0.011
(0.070) (0.174)
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.374 -0.222
(0.110) (0.148)
$FE_{ht}+FE_{cs}$ Yes Yes No No
$FE_{ht}+FE_{ct}+FE_{jt}$ No No Yes Yes
Observations 249,000 249,000 249,000 249,000
R-squared 0.04 0.05 0.32 0.32
Notes: This table reports estimates of the pass-through of statutory tariffs, $\Delta\ln(1+\tau_{cht})$, to duty-exclusive prices at the exporter–importer–product–year level, $\Delta\ln p_{ijht}$. Columns (2) and (4) include interactions between tariffs and lagged bilateral characteristics: supplier share ($s_{ijht-1}$), defined as supplier $i$’s share in buyer $j$’s imports of product $h$, and buyer share ($x_{ijht-1}$), defined as buyer $j$’s share in supplier $i$’s exports of product $h$. All regressions include the following controls: (i) $\ln\text{longevity}_{ijht}$, the number of years $i$ and $j$ have transacted in $h$; (ii) $\Delta\ln q_{i(-j)ht}$, exporter $i$’s sales of $h$ to U.S. buyers other than $j$; and (iii) $\Delta\ln p_{(-i)jht}$, the average price change charged by other suppliers of $h$ to buyer $j$, using lagged shares as weights. Columns (1)–(2) include product–year and exporter country–sector fixed effects ($FE_{ht}+FE_{cs}$), while Columns (3)–(4) include a more demanding set of fixed effects: product–year, importer–year, and exporter country–year ($FE_{ht}+FE_{ct}+FE_{jt}$). Standard errors are clustered at the HS8 product and exporter-country level. Arm’s length transactions are defined using LFTTD related party trade indicator. The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
@@ Appendix E Estimation Appendix
E.1 Downstream Demand Elasticity $\left(\nu\right)$
Consider a model where importer $j$ sells its output $q_{j}$ to downstream customers in different countries. A representative consumer in each country maximises utility by choosing a composite of domestic and imported goods. The sub-utility derived from the composite imported good will be given by a CES aggregation across imported varieties with a good-importer specific elasticity of substitution given by $\sigma_{g}$. Broda and Weinstein (2006) provide estimates of the elasticity $\sigma_{g}$ at the HS10 good $g$-level in U.S. import data. The plot below shows the distribution of these elasticities. We base the calibration of the elasticity $\nu$ in our model on these estimates. We consider a value of 4 for $\nu$, close to the mean value of 3.85, which we see as a conservative choice.

### v2

@@ D.4 Additional Results on Model Fit
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ -0.511 2.755 -0.278 2.816
(0.323) (0.461) (0.276) (0.617)
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$ -0.201 2.453 -0.130 2.703
(0.229) (0.242) (0.242) (0.349)
R-squared 0.10 0.36 0.36 0.49
@@
Notes: Panel A reports main effects of tariff changes on quantities. Panel B introduces interactions between tariff changes and lagged supplier share ($s_{ijht-1}$) and lagged buyer share ($x_{ijht-1}$). Columns (1) and (2) use baseline fixed effects ($FE_{ht}+FE_{cs}$), while Columns (3) and (4) use a more demanding specification with product–year, country–year, and buyer–year fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$). Standard errors are clustered at the HS8 product and exporter-country level. Observation counts are rounded to four significant digits per U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
@@
$\widehat{\Delta\ln q_{ijht}}$ 0.401 0.592 0.242 0.361 0.213 0.320 0.496 0.726
@@
$\widehat{\Delta\ln(p\cdot q)_{ijht}}$ 0.206 0.489 0.121 0.292 0.107 0.262 0.255 0.598
@@ D.5.1 Pass-Through Across Samples
Table D.11 examines how tariff pass-through estimates vary across alternative sample restrictions under baseline fixed effects (product–time and country–sector; Panel A) and more demanding specifications that additionally include country-time (instead of country-sector) and buyer–time fixed effects (Panel B). For reference, the first two columns replicate Columns (1) and (2) of Table 6. These estimates imply pass-through of roughly 82–88 percent in the data, closely aligned with the model’s predictions, with no statistically significant differences.
As the sample is progressively broadened–from including final goods (column 3), to incorporating single-buyer matches (column 4), and finally to the most inclusive specification, which adds related-party transactions, energy goods, and outliers (column 5)–estimated pass-through rises monotonically, reaching 94–96 percent.
These results indicate that sample composition plays an important role in measured pass-through. At the same time, when attention is restricted to stable buyer–supplier relationships, pass-through remains incomplete and below the near-100 percent estimates often reported in the literature. To further reconcile our findings with this evidence, we next aggregate the data to the product level and exploit monthly variation.
@@
Dependent variable: Model Baseline + Final +Suppliers + Energy/
$\Delta\ln p_{ijht}$ Goods w/ RPT/
<< 2 Buyers Outliers
@@
R-squared 0.48 0.31 0.26 0.14 0.12
Observations 249,000 249,000 473,000 1,490,000 1,768,000
Notes: This table reports tariff pass-through estimates to duty-exclusive CIF prices at the exporter–importer–product level. Panel A uses baseline fixed effects: product–time and country–sector ($FE_{ht}+FE_{cs}$). Panel B uses a more stringent specification: product–time, country–time, and buyer–time ($FE_{ht}+FE_{ct}+FE_{jt}$). Column (1) uses the model-predicted price change as the dependent variable. Column (2) uses the observed price change in the baseline sample and is identical to Column (1) of Table 3. Column (3) adds relationships in which the supplier trades with only one U.S. importer. Column (4) further expands the sample to include relationships that are either related parties, involve energy commodities, or exhibit extreme price levels or changes. Column (5) incorporates consumption goods, thus encompassing all consecutive exporter-importer-product combinations. Standard errors are clustered at the HS8 product and country level. Observation counts are rounded per U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
D.5.2 Product-Level Pass-Through Estimates
To complement our main analysis, we replicate a standard pass-through specification using data aggregated at the product-country-month level. We construct the data directly from the buyer–supplier–product triplets that form the basis of our firm-level regressions.
While much of the literature analyzes monthly price and tariff changes at the product–country level, such approaches reflect both intensive and extensive margin adjustments, including changes in trading partners or the entry and exit of relationships. In contrast, our aggregation focuses exclusively on consecutive transactions between the same buyer and supplier for a given product. This setup isolates price responses within ongoing relationships, capturing what is arguably the most direct expression of tariff pass-through at the micro level.
@@ D.5.2 Product-Level Pass-Through Estimates
We measure price changes at the HS10–source-country–month level using the same set of buyer–supplier–product links as in the baseline analysis, restricted to relationships with at least one transaction in both 2017 and 2018. For each relationship, we retain the month in which transactions occur in both years. We then aggregate values and quantities across these transactions to the product–source-country–month level, construct unit values, and compute month-to-month price changes.
Relative to Table D.11, this procedure introduces two changes: prices are aggregated across trading relationships to the product–source-country level, and price changes are measured at the monthly frequency rather than annually. The objective of Table D.12 is to evaluate how pass-through estimates change when prices are aggregated across firms within a product–source-country–month cell, while maintaining the same underlying set of buyer–supplier links observed in both 2017 and 2018.
Column (1) reports results for the baseline sample. The sample is then progressively broadened along the same margins as in Table D.11: Column (2) adds consumption goods, Column (3) allows suppliers serving only one buyer, Column (4) further includes related-party transactions, energy goods, and extreme price changes, and Column (5) additionally incorporates non-consecutive buyer–supplier relationships.
@@ Results
Across specifications, we continue to find incomplete tariff pass-through to U.S. import prices. When the aggregation is performed using the same set of repeated buyer–supplier relationships as in the baseline analysis, pass-through remains well below one, consistent with the match-level evidence. As the sample is progressively broadened to include consumption goods, single-buyer suppliers, related-party transactions, energy goods, and eventually non-consecutive relationships, estimated pass-through increases steadily, approaching the near-complete transmission typically found in the literature.
Differences between match-level and product-level pass-through therefore primarily reflect the set of relationships included in the data rather than aggregation itself. Incomplete pass-through remains a robust feature when attention is restricted to persistent buyer–supplier links.
@@
Table D.12: Product-Level Tariff Pass-Through Across Alternative Samples
Dependent variable: Baseline + Final + Suppliers + Energy/ + Non-Consecutive
$\Delta\ln p_{ijht}$ Goods w/ RPT/ Relationships
<< 2 Buyers Outliers
(1) (2) (3) (4) (5)
$\Delta\ln(1+\tau_{cht})$ -0.654 -0.353 -0.063 -0.054 -0.017
(0.245) (0.156) (0.073) (0.068) (0.068)
$FE_{ht}+FE_{cs}$ Yes Yes Yes Yes Yes
R-squared 0.23 0.22 0.14 0.13 0.11
Observations 124,000 213,000 665,000 805,000 1,071,000
@@
Notes: This table reports product-level pass-through estimates. Product-level observations are constructed from firm-buyer-HS relationships that transact in both years t and t-1. Values and quantities are aggregated to the product–source-country–month level, unit values are constructed at that frequency, and month-to-month changes in log CIF prices are computed. Column (1) reports the baseline sample, restricted to repeated arm’s-length relationships in intermediate goods among multi-buyer suppliers. Column (1) aggregates the baseline sample used in Table 3 to the product–source-country level. Column (2) broadens the sample to include consumption goods (all BEC categories). Column (3) further adds suppliers that trade with only one U.S. importer. Column (4) further includes related-party transactions, energy commodities, and observations with extreme price levels or changes. Column (5) removes the consecutive-year restriction and includes all exporter–importer–product combinations. Standard errors are clustered at the HS8 product and country level. Observation counts are rounded in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
@@ Appendix E Estimation Appendix
E.1 Calibration of Downstream Demand Elasticity $\left(\nu\right)$
Consider an extension to our model in which importer $j$ sells its output $q_{j}$ to downstream consumers in different countries. In each country, a representative consumer allocates expenditure between domestic and imported goods, and the sub-utility over imported varieties is CES with elasticity of substitution $\sigma_{g}$ at the HS10 good level. Broda2006 (BW) estimate these import demand elasticities for the United States.
In our framework, the parameter $\nu$ governs the own-price elasticity of demand faced by importer $j$’s output in the downstream market. We interpret BW’s $\sigma_{g}$ estimates as measuring this variety-level demand elasticity. Accordingly, we calibrate $\nu$ using their estimates. The distribution of $\sigma_{g}$ is shown in Figure E.1. Given a mean value of 3.85 and a median of 2.8, we set $\nu=4$, which we view as a conservative choice.